The numerator of a fraction is 7 less than the denominator. If 3 is added to the numerator and 9 is subtracted from the denominator the resulting fraction is equal to 3/2. Find the original fraction.
I tried the following, but am not finding the right answer:
"Let x= denominator"
"Let x-7= numerator"
(X-9)+(x-7)+9=3/2
2(x-9+x-7+9)= 3/2 (2/1)
2x-18+2x-14+18=3
4x-14=3
4x=17
X=17/4

Answer by richwmiller(17219)

x-7/x
(x-7+3)/(x-9)=3/2
(x-4)/(x-9)=3/2
2*(x-4)=3*(x-9)
2x-8=3x-27
x=19
12/19 original fraction
check
12/19
15/10=3/2
ok